OpenStudy (anonymous):

A particle is moving with the given data. Find the position of the particle. a(t)=sin(t)+3cos(t) s(0)=0 v(0)=2

OpenStudy (unklerhaukus):

a(t) = dx^2/dt^2 integrate twice

OpenStudy (anonymous):

\[v(t) = \int\limits_{}^{}a(t) + C\] plug in v(0) = 2 to solve for C to get the real v(t) for this case integrate and repeat to find the function s(t)

OpenStudy (anonymous):

i did not get it. could you explain it more?

OpenStudy (anonymous):

s(t)= - sin(t) - 3 cos (t) + 3(t) + 3 i got this answer but i want to make sure is it correct or not